OpenStudy (anonymous):

A farmer with 2070 meters of fencing wants to enclose a rectangular plot that borders on a straight highway. If the farmer does not fence the side along the highway, what is the largest area that can be enclosed? Give your answer correct to the nearest square meter.

OpenStudy (amistre64):

2070/3 per side would make a square; and a square holds the largest area

OpenStudy (anonymous):

sooo 2070/3=690 690*690=476100 but thats not right .. sorry i suck at math

OpenStudy (amistre64):

476,100 is the right math to it

OpenStudy (amistre64):

lets work it out the long way ..... i spose :)

OpenStudy (anonymous):

ok:)

OpenStudy (amistre64):

Area = xy P = 2x + y = 2070; y = 2070 -2x Area = x(2070 -2x) = 2070x -2x^2 Area' = 2070 -4x, =0 at x = 2070/4

OpenStudy (anonymous):

its for a webassign (online quiz) and i put that answer in but it keeps saying its wrong :(

OpenStudy (amistre64):

y = 2070 - 2(2070)/4 = 2070 - 2070/2 = 2070 - 1035 = 1035 try; 2070(1035)/4

OpenStudy (anonymous):

ok ill try it ! :)

OpenStudy (amistre64):

535612.5 is the max area then

OpenStudy (anonymous):

yesss :) that was correct thanks sooooo much :)

OpenStudy (anonymous):

can you help me with one more?